# One-step Generative Diffusion for Realistic Extreme Image Rescaling

### v2

@@ 3.3 One-Step Perception Enhancement
where $\hat{z}_{0}$ represents the perceptually enhanced latent features, $\bar{\alpha}_{t}$ is the noise scheduler of the forward diffusion process of the SD model, and $t$ represents the time-step associated with the noise intensity. Finally, $\hat{z}_{0}$ can be transformed to the pixel domain through the pre-trained VAE decoder $\mathcal{D}$ to obtain the upscaled image $\hat{x}$:
$\hat{x}=\mathcal{D}\left(\hat{z}_{0}\right)$ (10)
Then, we can adopt the loss function between $\hat{x}$ and $x$ to optimize the U-Net using LoRA layers.
It should be noted that one of the key designs is the selection of the time-step $t$. Previous studies [35] have shown that the time-step in diffusion models could determine the trade-off between realism and fidelity. A larger $t$ results in stronger diversity and greater realism, but at the cost of fidelity. However, we have observed that the quality of the rescaled latent features exhibit non-uniform characteristics, as shown in Fig. 4. After rescaling, the low-frequency regions of an image typically experience less degradation, while the degradation in high-frequency areas is more conspicuous, necessitating a higher generative capability for restoration. Thus, we propose to dynamically predict the time-step $t$ based on the quality of the rescaled latent features. We propose to use a simple yet effective network named time-step prediction module (TPM) to regress $t$ based on the rescaled latent features:
$t=\mathrm{TPM}\left(\hat{z}\right)$ (11)
@@
Subsequently, the predicted time-step $t$ is used to perform the single denoising step. During the training phase, we compute a single time-step for each input image patch. In the inference phase, we adopt a tiled inference approach, allowing for predicting different time-steps for various regions of a single image. However, the fixed scheduler used in equation (9) is not differentiable with respect to $t$, which prevents gradients from being back-propagated to the TPM. Considering that given $\hat{z}$, $\epsilon$ and $t$, equation (9) essentially represents a linear transformation, which can be learned by a neural network. However, if we simply use a randomly initialized neural network to simulate this process, it would cause the output to deviate significantly from the latent space of the pre-trained VAE, leading to training instability. Therefore, we propose a hybrid refiner module consisting of both a fixed scheduler and a learnable scheduler. The former uses equation (9) to perform the denoising step, while the latter employs a neural network to simulate the denoising scheduler. To ensure training stability, the final layer of the learnable scheduler is a zero-initialized convolution, allowing it to gradually correct the results of the fixed scheduler as training progresses. This process can be described as follows:
@@
Here $t_{0}$ represents the fixed pre-set time-step.
Figure 5: The structure of the refiner module. It combines both the fixed scheduler and learnable scheduler to perform the denoising process. The two schedulers are connected by a zero-convolution layer to stabilize the training.
@@
DIV2K EDSR 24.64 22.13 0.6572 0.6056 0.5376 0.6219 0.3162 0.4184 34.15 23.28 0.3397 0.4193
RCAN 24.34 22.06 0.6479 0.6034 0.5485 0.6307 0.3269 0.4349 31.57 23.18 0.3408 0.4091
ESRGAN 23.15 21.15 0.5946 0.5808 0.4478 0.5788 0.2378 0.4237 59.80 43.40 0.6161 0.3725
@@
Ours 23.77 22.10 0.6675 0.5839 0.3026 0.4281 0.0824 0.1597 66.31 65.51 0.6443 0.6521
Urban100 EDSR 19.93 18.08 0.5020 0.4447 0.5827 0.6839 0.3701 0.4734 41.97 24.80 0.3439 0.4173
RCAN 19.96 18.14 0.5017 0.4452 0.5859 0.6880 0.3745 0.4808 40.31 24.46 0.3561 0.4130
ESRGAN 19.36 17.68 0.4835 0.4344 0.4741 0.6142 0.2554 0.4288 65.82 57.02 0.5849 0.3860
@@
Ours 21.34 18.77 0.6464 0.4655 0.2601 0.4490 0.1107 0.2265 72.21 71.03 0.6773 0.6643
CLIC2020 EDSR 26.33 23.92 0.7408 0.7062 0.5109 0.5742 0.3057 0.4024 33.17 21.77 0.3392 0.4307
RCAN 26.37 24.04 0.7407 0.7072 0.5133 0.5777 0.3126 0.4151 31.59 22.07 0.3469 0.4180
ESRGAN 25.43 23.48 0.7018 0.6991 0.4213 0.5347 0.2209 0.4014 57.80 40.98 0.6107 0.3537
Ours 25.81 24.16 0.7497 0.6887 0.2777 0.3968 0.0754 0.1459 65.84 64.38 0.6198 0.6586
DIV8K EDSR 26.56 24.48 0.6776 0.6472 0.5437 0.6048 0.3358 0.4325 30.83 22.30 0.3701 0.4686
RCAN 26.63 24.57 0.6778 0.6481 0.5470 0.6085 0.3430 0.4449 29.93 22.52 0.3796 0.4502
ESRGAN 25.49 23.78 0.6247 0.6350 0.4628 0.5610 0.2518 0.4268 53.49 39.11 0.6401 0.4078
@@
Ours 25.93 24.15 0.7107 0.6302 0.3198 0.4376 0.0927 0.1798 57.75 58.89 0.6598 0.6461
Figure 6: Qualitative comparisons of 16$\times$ (the 2nd and 3rd rows) and 32$\times$ (the 4th and 5th rows) rescaling results. The first row represents the original high-resolution images. Our method achieves higher semantic accuracy, such as more recognizable faces, clearer text, and more realistic structures.
Figure 8: Visual comparisons of 32$\times$ rescaling methods on the DIV8K dataset
In this part, we fine-tune the denoising U-Net network using LoRA, while also jointly training the TPM and refiner module. To achieve a balance between fidelity and realism, we supervise the reconstruction results $\hat{x}$ using pixel loss and perceptual loss as follows:
$\begin{split}\mathcal{L}_{\mathrm{enh.}}=&\underbrace{\left\|\hat{x}-x\right\|_{1}}_{\text{pixel loss}}+\\ &\underbrace{\lambda_{\mathrm{lpips}}\mathcal{L}_{\mathrm{lpips}}\left(\hat{x},x\right)+\lambda_{\mathrm{dists}}\mathcal{L}_{\mathrm{dists}}\left(\hat{x},x\right)}_{\text{perceptual loss}}\end{split}$ (13)
@@ 4.1 Experimental Setup
We train the proposed OSIRDiff using the DF2K dataset [1, 19]. Considering that traditional benchmarks such as Set5, Set14, and BSD100 often have lower quality and resolution [5], they are not suitable for evaluating extreme rescaling methods. Therefore, we utilize four high-resolution datasets: DIV2K, Urban100 [7], CLIC2020 [32], and DIV8K [5] to evaluate models been compared. We adopt PSNR and SSIM [37] for quantitative evaluation. For perceptual quality evaluation, reference-based metrics LPIPS [49] and DISTS [3] are used. For non-reference image quality assessment, we use metrics including MUSIQ [13] and CLIPIQA [33].
@@ 4.1.2 Training Details
OSIRDiff is built based on the SD 2.1-base model, and the training is divided into three stages. First, we train the DFRM using equation (8). Then we adopt equation (13) to jointly train the LoRA layers, the TPM and the refiner module. Finally, we use a smaller learning rate to jointly fine-tune DFRM, LoRA layers, TPM and refiner module using a combination of equations (8) and (13). More training details are provided in the supplementary material.
@@ 4.2 Results and Comparison with SOTA
We compare the proposed OSIRDiff with three kinds of methods: (1) bicubic downscaling and SR, i.e., EDSR [19], RCAN [50] and ESRGAN [36]; (2) flow-based rescaling models, i.e., IRN [41] and HCFlow [18]; (3) prior-based extreme rescaling method VQIR [39]. Some of them are proposed for 4$\times$ rescaling factors and we extended them to 16$\times$ or 32$\times$. To ensure a fair comparison, all methods are trained until convergence on the DF2K dataset following their original configurations.
@@ 4.2.2 Qualitative Results
In Fig. 6, we present visual comparisons at 16$\times$ and 32$\times$ rescaling factors. Due to the ill-posed nature of extreme rescaling, pixel-oriented optimization methods such as EDSR, RCAN, IRN, and HCFlow produce overly-smoothed results. Although ESRGAN incorporates perceptual optimization it still cannot generate adequate details. VQIR leverages the pre-trained VQGAN as a prior, allowing it to generate rich details and sharp edges. However, it struggles with faces and more complex structures. Thanks to the powerful prior provided by the pre-trained SD and the rich semantic information contained in the latent space of the pre-trained VAE, our method effectively addresses these issues. It not only generates sufficiently rich textures but also reconstructs accuracy semantics and regular structures, such as faces, text and architectural structures.
The quantitative experiments demonstrate that our model exhibits superior performance on the Urban100 [7] dataset, confirming that our method produces more accurate structural features. In this section, we provide a visual comparison of 16$\times$ rescaling results on the Urban100 dataset, as shown in Fig. 7. Traditional regression models, such as EDSR [19], RCAN [50], IRN [41], and HCFlow [18], tend to generate overly-smooth results. GAN-based methods, including ESRGAN [36] and VQIR [39], generate more details, but their edge information lacks accuracy and regularity. In contrast, our approach, by leveraging the prior knowledge encapsulated in the SD model, is able to reconstruct more accurate and rich edge details. In Fig. 8, we present a visual comparison of 32$\times$ rescaling results on the ultra-high-resolution dataset, DIV8K [5]. It is evident that our method is capable of recovering images with accurate semantics, even at extreme upscaling factors.
@@ 4.3 Ablation Study
We conduct ablation experiments to validate the effectiveness of each component. All ablations are trained on the DF2K dataset and tested on the validation set of the DIV2K.
@@ 4.3.1 Effectiveness of Feature Rescaling
Table 2: Ablation study of rescaling space and SD prior.
Latent space Pixel space SD prior LPIPS\ $\downarrow$ DISTS\ $\downarrow$
$\times$ $\checkmark$ $\checkmark$ 0.3630 0.1154
$\checkmark$ $\times$ $\times$ 0.4675 0.3109
$\checkmark$ $\times$ $\checkmark$ 0.3026 0.0824
@@ 4.3.3 Effectiveness of DFRM Design
4.3.4 Effectiveness of Time Alignment
To validate the effectiveness of our proposed time-step alignment strategy, we retrain our model using fixed time-steps. This involves removing the TPM and refiner modules and employing a manually selected time-step to perform a single denoising step based on equation (9). Following the setting of previous work [40], we select fixed time-steps 1 and 999 for comparison. As shown in Fig. 14, when the time-step is set to a smaller value, the model achieves higher fidelity but lower perceptual quality, and the situation is reversed when a larger time-step is chosen. In contrast, our time-step alignment mechanism can dynamically determine the level of generative capability based on the image content, thereby achieving optimal performance in both fidelity and perceptual quality.
5 Discussions
5.1 Comparison with JPEG Compression
Our model can be combined with existing lossless image compression methods to achieve extreme compression ratios. To validate this, we use the lossless compression image format PNG to compress LR images and compare its performance with JPEG compression at various bit-rates. As shown in Fig. 15, at the same bits per pixel (bpp), the image reconstruction quality of our model significantly surpasses that of the JPEG compression. These comparisons demonstrate the potential of our model for image compression.

### v3

@@ 3.3 Denoising Guided Perception Enhancement
where $\hat{z}_{0}$ represents the perceptually enhanced latent features, $\bar{\alpha}_{t}$ is the noise scheduler of the forward diffusion process of the SD model, and $t$ represents the time-step associated with the noise intensity. Finally, $\hat{z}_{0}$ can be transformed to the pixel domain through the pre-trained VAE decoder $\mathcal{D}$ to obtain the upscaled image $\hat{x}=\mathcal{D}\left(\hat{z}_{0}\right)$. Then, we can adopt the loss function between $\hat{x}$ and $x$ to optimize the U-Net using LoRA layers.
However, the key problem of the aforementioned approaches lies in the selection of time-step $t$. We can regard the rescaling process as the noise-adding process of diffusion model, while Eq. (9) represents the denoising process. For different rescaling factors and different images, the ‘noise’ introduced by the rescaling operation varies in intensity. In Fig. 3, we present the global rescaling error of an image under different rescaling factors, as well as the local rescaling error of different patches within the image. Additionally, we plot the error curve of the forward noise-adding process of the SD model. It can be observed that as the rescaling factor and image content change, rescaling MSE also varies and corresponds to the diffusion MSE at different time-steps. Thus, we propose to dynamically predict the time-step $t$ based on the quality of the rescaled latent features $\hat{z}$. A simple yet effective network named time-step prediction module (TPM), whose structure is presented in Fig. 1, is used to predict $t$ based on the rescaled latent features: $t=\mathrm{TPM}\left(\hat{z}\right)$. During the training phase, we compute a single time-step for each input image patch. In the inference phase, we adopt a tiled inference approach (which will be elaborated in detail in the supplementary materials), allowing for predicting different time-steps for various regions of a high-resolution input image.
Figure 3: We compare the MSE introduced by the rescaling operation with that caused by the forward diffusion process. The rescaling degradation severity is correlated with both the image content and the rescaling scale. Therefore, the model requires adaptive time-step to align the rescaling MSE with the diffusion MSE.
@@
Subsequently, the predicted time-step $t$ is used to perform a single denoising step. However, the vanilla SD time scheduler as shown in Eq. (9) is not differentiable with respect to $t$, which prevents gradients from being back-propagated to the TPM. Considering that given $\hat{z}$, $\epsilon$ and $t$, Eq. (9) essentially represents a linear transformation, which can be learned by a neural network. Therefore, we propose a hybrid time scheduler module consisting of a fixed scheduler and a learnable scheduler, whose structure is presented in Fig. 4. The former uses Eq. (9) to perform the denoising step, while the latter employs a neural network to simulate the denoising scheduler. To ensure training stability, the final layer of the learnable scheduler is a zero-initialized convolution, allowing it to gradually correct the results of the fixed scheduler as training progresses. This process can be described as follows:
@@
where $t_{0}$ is the fixed pre-set time-step and $\mathcal{S}_{\mathrm{fixed}}$ denotes the vanilla SD time scheduler as shown in Eq. (9).
Figure 4: The structure of the time scheduler module. It combines both the fixed scheduler and learnable scheduler to perform the denoising process. The two schedulers are connected by a zero-convolution layer to stabilize the training.
@@
DIV2K ESRGAN 23.15 19.94 0.5946 0.5689 0.4478 0.5892 0.2378 0.4230 59.80 43.35 0.6161 0.3734
StableSR 22.02 19.92 0.5983 0.5588 0.4980 0.5736 0.2078 0.2603 51.28 48.94 0.3780 0.3203
S3Diff 20.22 17.81 0.5491 0.4601 0.4033 0.4895 0.1309 0.1810 64.37 67.92 0.6228 0.6991
@@
Ours 23.98 22.18 0.6736 0.5869 0.2979 0.4221 0.0886 0.1684 66.56 69.12 0.7189 0.7204
CLIC2020 ESRGAN 25.43 22.47 0.7018 0.6804 0.4213 0.5429 0.2209 0.4011 57.80 41.28 0.6107 0.3558
StableSR 24.11 22.24 0.7052 0.6726 0.4625 0.5288 0.1860 0.2282 52.69 49.80 0.3779 0.3325
S3Diff 22.04 19.87 0.6615 0.5711 0.3968 0.4829 0.1336 0.1812 63.07 66.40 0.6001 0.6838
Ours 26.04 24.26 0.7542 0.6903 0.2724 0.3911 0.0792 0.1513 65.35 69.10 0.6815 0.7271
Urban100 ESRGAN 19.36 17.68 0.4835 0.4344 0.4741 0.6142 0.2554 0.4288 65.82 57.02 0.5849 0.3860
StableSR 19.93 18.08 0.5020 0.4447 0.5827 0.6839 0.3701 0.4734 41.97 24.80 0.3439 0.4173
S3Diff 17.60 15.74 0.4368 0.3419 0.4204 0.5282 0.1917 0.2573 71.01 69.56 0.6658 0.6776
@@
Ours 21.46 18.84 0.6495 0.4685 0.2577 0.4481 0.1154 0.2386 72.22 72.88 0.7155 0.6795
DIV8K ESRGAN 25.49 23.78 0.6247 0.6350 0.4628 0.5610 0.2518 0.4268 53.49 39.11 0.6401 0.4078
StableSR 25.90 23.55 0.6602 0.6258 0.4844 0.5497 0.2134 0.2623 47.15 46.97 0.4017 0.3557
S3Diff 22.20 20.19 0.5903 0.5063 0.4163 0.4954 0.1524 0.1997 59.98 62.58 0.6351 0.6901
@@
Ours 26.16 24.37 0.7163 0.6337 0.3117 0.4358 0.0979 0.1888 58.54 63.34 0.6836 0.7229
Finally, we fine-tune the denoising U-Net network and the VAE decoder using LoRA, while also jointly training the TPM and time scheduler module. To achieve a balance between fidelity and realism, we supervise the reconstruction results $\hat{x}$ using pixel loss and perceptual loss as follows:
$\mathcal{L}_{\mathrm{enh.}}=$ $\left\|\hat{x}-x\right\|_{1}+$ (11)
$\lambda_{\mathrm{pec.}}\left(\mathcal{L}_{\mathrm{lpips}}\left(\hat{x},x\right)+\mathcal{L}_{\mathrm{dists}}\left(\hat{x},x\right)\right)$
@@ 4.1 Experimental Setup
We train the proposed TADM using the DF2K dataset [1, 18]. Considering that traditional benchmarks such as Set5, Set14, and BSD100 often have lower quality and resolution [5], they are not suitable for evaluating extreme rescaling methods. Therefore, we utilize four high-resolution datasets: DIV2K, CLIC2020 [25], Urban100 [6], and DIV8K [5] to evaluate models been compared. We adopt PSNR and SSIM [31] for quantitative evaluation. For perceptual quality evaluation, reference-based metrics LPIPS [45] and DISTS [2] are used. For non-reference image quality assessment, we use metrics including MUSIQ [11] and CLIPIQA [26].
@@ 4.1.2 Training Details
TADM is built based on the SD 2.1-base model, and the training is divided into three stages. First, we train the DFRM using Eq. (8). Then we adopt Eq. (11) to jointly train the LoRA layers, the TPM and the time scheduler module. Finally, we use a smaller learning rate to jointly fine-tune DFRM, LoRA layers, TPM and scheduler module using a combination of Eq. (8) and (11). More training details are provided in the supplementary material.
@@ 4.2 Results and Comparison with SOTA
We compare the proposed TADM with three kinds of methods: (1) bicubic downscaling and SR, i.e., ESRGAN [29], StableSR [27] (based on multi-step diffusion) and S3Diff (based on single-step diffusion)[42]; (2) flow-based rescaling models, i.e., IRN [36] and HCFlow [17]; (3) prior-based extreme rescaling method VQIR [33]. Some of them are proposed for 4$\times$ rescaling factors and we extended them to 16$\times$ or 32$\times$. To ensure a fair comparison, all methods are trained until convergence on the DF2K dataset following their original configurations.
@@ 4.2.1 Quantitative Results
Figure 5: Qualitative comparisons of 16$\times$ (the 1st and 2nd rows) and 32$\times$ (the 3rd and 4th rows) rescaling results. Our method achieves higher semantic accuracy, such as more clearer text, more recognizable faces, and more realistic structures.
Figure 7: Visual comparisons of 32$\times$ rescaling methods on the DIV8K dataset. Our approach can still reconstruct correct semantic information even in such extreme scenarios.
@@ 4.2.2 Qualitative Results
In Fig. 5, we present visual comparisons at 16$\times$ and 32$\times$ rescaling factors. Due to the ill-posed nature of extreme rescaling, pixel-oriented optimization methods such as IRN, and HCFlow produce overly-smoothed results. Although ESRGAN incorporates perceptual optimization it still cannot generate adequate details. VQIR leverages the pre-trained VQGAN as a prior, allowing it to generate rich details and sharp edges. However, it struggles with faces and more complex structures. Although S3Diff can generate sharp edges and rich details, its reconstructed results have low fidelity and exhibit significant artifacts and color distortions. Thanks to the powerful prior provided by the pre-trained SD and the carefully designed architecture,, our method effectively addresses these issues. It not only generates sufficiently rich textures but also reconstructs accuracy semantics and regular structures, such as text, faces, vehicles, and leaves.
The quantitative experiments demonstrate that our model exhibits superior performance on the Urban100 [6] dataset, confirming that our method produces more accurate structural features. In this section, we provide a visual comparison of 16$\times$ rescaling results on the Urban100 dataset, as shown in Fig. 6. Traditional regression models, such as IRN [36] and HCFlow [17], tend to generate overly-smooth results. GAN-based methods, including ESRGAN [29] and VQIR [33], generate more details, but their edge information lacks accuracy and regularity. For diffusion-based super-resolution methods, such as S3Diff [42], they are capable of generating sharp edges. However, due to the lack of information about downscaling, their results often exhibit lower fidelity, which manifests in distorted structures. In contrast, our approach, by leveraging the prior knowledge encapsulated in the SD model, is able to reconstruct more accurate and rich edge details. In Fig. 7, we present a visual comparison of 32$\times$ rescaling results on the ultra-high-resolution dataset, DIV8K. It is evident that our method is capable of recovering images with accurate semantics, even at extreme rescaling factors. For instance, our method is capable of accurately reconstructing electronic devices, doors, and chairs, while retaining rich details.
@@ 4.3 Ablation Study
We conduct ablation experiments to validate the effectiveness of each component. All ablations are trained on the DF2K dataset and tested on the validation set of the DIV2K for 16$\times$ rescaling task.
Table 2: Ablation study of rescaling space and SD prior.
Latent space Pixel space SD prior LPIPS\ $\downarrow$ DISTS\ $\downarrow$
$\times$ $\checkmark$ $\checkmark$ 0.3630 0.1154
$\checkmark$ $\times$ $\times$ 0.4675 0.3109
$\checkmark$ $\times$ $\checkmark$ 0.2979 0.0886
@@ 4.3.3 Effectiveness of DFRM Design
4.3.4 Effectiveness of Time Alignment
To validate the effectiveness of our proposed time-step alignment strategy, we retrain our model using fixed time-steps. This involves removing the TPM and time scheduler modules and employing a manually selected time-step to perform a single denoising step based on Eq. (9). Following the setting of previous work [34], we select fixed time-steps 1 and 999 for comparison. As shown in Fig. 13, when the time-step is set to a smaller value, the model achieves higher fidelity but lower perceptual quality, and the situation is reversed when a larger time-step is chosen. In contrast, our time-step alignment mechanism can dynamically determine the level of generative capability based on the image content, thereby achieving optimal performance in both fidelity and perceptual quality.
@@ 4.3.4 Effectiveness of Time Alignment
Figure 14: Illustration of the tiled inference process and the predicted time-step map.
In Fig. 14, we present a schematic illustration of the tiled inference process along with the predicted time-step map. It can be observed that the time alignment mechanism effectively captures more complex regions in high-resolution images and assigns larger time steps to these areas.
4.4 Discussions
4.4.1 Details about Tiled Inference
When TADM is employed for rescaling ultra-high-resolution images, it often necessitates dividing the input image into multiple patches for separate processing. In the main paper, this is referred to as the tiled inference strategy. In Algorithm 1, we elaborate on the process of tiled inference. Unlike works such as StableSR [27], our tiled inference algorithm can predict different time steps t for each image patch, thereby achieving dynamic allocation of generative capacity.
Algorithm 1 Tiled Inference Process
1:Input image $x$, latent encoder $\mathcal{E}$, decoupled feature rescaling module $\mathrm{DFRM}$, denoising U-Net $\epsilon_{\theta}$, time prediction module $\mathrm{TPM}$, time scheduler $\mathrm{TS}$, latent decoder $\mathcal{D}$, patch size $p$, stride length $s$
2:Rescaled image $\hat{x}$, LR image $y$
3:$z=\mathcal{E}(x)$triangleright latent encoding
4:$\hat{z},y=\mathrm{DFRM}(z)$triangleright latent rescaling
5:$[\hat{z}_{i}]=\mathrm{ToPatch}(\hat{z},p,s)$triangleright split $\hat{z}$ to patches
6:$z_{0}\_{list}=[]$triangleright initialize list of $z_{0}$
7:for $\hat{z}_{i}$ in $[\hat{z}_{i}]$ do
8:     $t_{i}=\mathrm{TPM}(\hat{z}_{i})$triangleright time-step prediction
9:     $\epsilon_{i}=\epsilon_{\theta}(\mathrm{patch},t_{i})$triangleright noise prediction
10:     $z_{0}^{i}=\mathrm{TS}(\hat{z}_{i},\epsilon,t_{i})$triangleright denoising by time scheduler
11:     $z_{0}\_{list}.\mathrm{append}(z_{0}^{i})$
12:end for
13:$z_{0}=\mathrm{Merge}(z_{0}\_{list})$triangleright merge patches of $z_{0}$
14:$\hat{x}=\mathcal{D}(z_{0})$triangleright latent decoding
15:return $\hat{x},y$triangleright output
The aforementioned algorithm encompasses two hyperparameters: the patch size and the stride length, both defined in the latent space. If the patch size is set too large, although inference efficiency may improve, the prediction of time steps would become overly sparse, leading to performance degradation. Conversely, if the patch size is set too small, the computational load will be significantly higher, and there might be a misalignment between the inference size and the pre-training image size of Stable Diffusion.
Therefore, to achieve a trade-off between performance and inference efficiency, we conduct experiments on the patch size, as shown in Fig. 15. It can be observed that the model achieves optimal performance when the patch size is between 60 and 120. Consequently, we select 96 as the patch size in our work. For the stride length, we adopt the value from previous works [27] and directly set it to 64.
In Fig. 16, we present the time step predictions when employing different patch sizes during the tiled inference process. It can be observed that smaller patch sizes allow for finer-grained time step predictions. However, setting the size too small leads to misalignment with the pre-training image size of Stable Diffusion, resulting in increased computational load and performance degradation. In contrast, our selected patch size of 96 achieves a balance between performance and computational load, while also enabling relatively accurate prediction of the time step mask.
Figure 15: Ablation study about patch size.
Figure 16: The time step predictions at different patch sizes.
4.4.2 Application to Image Compression
Our model can be combined with existing lossless image compression methods to achieve extreme compression ratios. To validate this, we use PNG to compress LR images and compare it with JPEG at various bit-rates. As shown in Fig. 17, at the same bits per pixel (bpp), the image reconstruction quality of our model significantly surpasses that of the JPEG compression. These comparisons demonstrate the potential of our model for image compression.